# Multilevel Control Functional

### v3

@@ 4 Experimental Results
To compare MCLF using quasi-Monte Carlo points (QMC), Latin hypercube sampling (LHS), and i.i.d points (IID) with MLMC and CF using i.i.d points, we repeat the experiment 100 times. Multilevel methods in this example utilize the optimal sample size for MLMC. Details about the sample size, evaluation cost at each level, and other relevant information can be found in Appendix D.3. As shown in Figure˜4, under the same evaluation cost constraint, MLCF outperforms MLMC and CF. Figure˜4 also shows that experimental designs can improve the performance of MLCF.
@@
We now consider to perform Bayesian inference for the Lotka-Volterra system (Lotka, 1925, 1927; Volterra, 1927), which is also known as the predator-prey model. The model usually uses a system of differential equations:
@@
to describe the interaction between a predator and its prey in an ecosystem. $u_{1}(t)$ and $u_{2}(t)$ are the prey population and the predator population at time $t\in[0,s]$, for some $s\in\mathbb{R}_{+}$. The initial conditions of the system are $u_{1}(0)=x_{5}$ and $u_{2}(0)=x_{6}$. The observations $u_{1}(t_{i})$ and $u_{2}(t_{i})$ obtained exhibit log-normal noise with independent standard deviation $x_{7}$ and $x_{8}$ respectively, for all $i\in\{1,\ldots,m\}$. We can re-parameterize $x$ as in (Sun et al., 2023b, a) such that the re-parameterized model has parameters $\tilde{x}\in\mathbb{R}^{8}$. With the Gaussian distribution priors we assign on $\tilde{x}$ and the observations, we can construct the posterior distribution of $\tilde{x}$. The quantity of interest $\Pi[f]$ is the posterior expectation of the average prey population over the time period between 0 and $s$, i.e. $\Pi[f]=\int_{\mathbb{R}^{8}}f(\tilde{x})\pi(\tilde{x})d\tilde{x}$, where $\pi$ is the posterior probability distribution of $\tilde{x}$ and f(x~)f($\tilde${x}) is the average prey population between 0 and $s$ with the model parameter $\tilde{x}$. $f(\tilde{x})=s^{-1}\int_{0}^{s}u_{1}(t)dt$ is approximated with $(s)^{-1}h\sum^{s/h}_{i=1}u_{1}(t_{i})$, where $h$ is the step size and each $u_{1}(t_{i})$ is obtained by solving the differential equations numerically. The real-world dataset (Hewitt, 1921) consisting of the population of snowshoe hares (prey) and Canadian lynxes (predators) is used as observations for our study. With the real-world observations, we conduct Bayesian inference and use a MCMC sampler (no-U-turn sampler) in Stan (Carpenter et al., 2017) to obtain samples.
@@
We further extend the proposed MLCF estimators for variational inference. In particular, we applied a Bayesian neural network regression model to the UCI wine-quality-red dataset as (Fujisawa and Sato, 2021). The Bayesian neural network consists of a $15$-unit or a $20$-unit hidden layer with ReLU activations. For each weight $w_{i}$, we set its prior as Gaussian $w_{i}\sim N(\mu_{w},\sigma_{w})$, and the response variable $y\sim N(\phi(x,\{w_{i}\}_{i=1}^{d}),\sigma)$. The model then have a posterior of $d=392$ (when the number of hidden units is $15$) or $d=522$ (when the number of hidden units is $20$) and was applied to a dataset sub-sampled from the wine-quality-red dataset. The posterior is approximated by deploying a diagonal Gaussian distribution as the variational distribution. We compare the MLCFRG with MLMCRG, MLMC (with the number of levels being $2$) and MC estimators. For MLCFRG, we use preconditioned squared-exponential kernels (Oates et al., 2017); see details in Appendix C.1. The results, the average training ELBO/test log-likelihood (averaged over 10 runs) and associated standard deviations, are illustrated in Figure˜6 and Figure˜6. For each run, the networks are initialized under identical settings. See Appendix D.5 for extra details and experiments. Empirical results show that the MLCFRG estimators tend to outperform the other estimators as they have shown a faster convergence rate towards the optimal training ELBO and testing log-likelihood over number of iterations or equivalently the number of training data.
@@ A.2 Stein kernels
$k_{0}(x,x^{\prime})$ $=\nabla_{x}\cdot\nabla_{x^{\prime}}k(x,x^{\prime})+\nabla_{x}\log\pi(x)\cdot\nabla_{x^{\prime}}k(x,x^{\prime})+\nabla_{x^{\prime}}\log\pi(x^{\prime})\cdot\nabla_{x}k(x,y)$
@@ Appendix B Proofs of Theoretical Results
due to the property of the Stein kernel $k^{l}_{0}$ that the Stein kernel $k^{l}_{0}$ satisfies $\int_{\mathcal{X}}k^{l}_{0}(x,x^{\prime})\pi(x)dx=0$ for all $x\in\mathcal{X}$ [Oates et al., 2017, 2019]. Then, we have
@@
Following the proof of Theorem 1 of [Oates et al., 2019] or Theorem 11.13 of [Wendland, 2004], under assumptions A1-7, there exists $r^{*}_{l}>0$ and $h^{*}_{l}>0$, for $h_{l}<h^{*}_{l}$,
@@ C.1 Kernels and Their Derivatives
A preconditioned squared-exponential kernel [Oates et al., 2017] is:
@@ C.2 Hyper-parameter Selection
We propose to select kernel hyperparameters $\nu$ through a marginal likelihood objective by noticing the equivalence between the optimal Stein kernel-based control variates on the objectives in [Oates et al., 2017, Sun et al., 2023b] and the posterior mean of a zero-mean Gaussian process model with covariance matrix $k_{0}(x,y)$. See [Oates et al., 2017] for a discussion in the scalar-valued control variates case; see [Sun et al., 2023b] for a discussion in the vector-valued control variates case. Denotes all kernel-hyperparameters by $\nu:=\{\nu_{1},\ldots,\nu_{L}\}$, and we maximize the sum of the marginal likelihood,
@@ D.3 Experimental Details for the ODE example
We follow the settings described in [Giles, 2015, Li et al., 2023]. For completeness, we provide details of the solver (finite difference approximation). The boundary-value ordinary differential equation (ODE) with random coefficient and random forcing is given by:
@@
For the ODE example, we have evaluation cost at each level $C=(C_{0},C_{1},C_{2})=(1.22,3.57,11.89)$ (all measured in $10^{-3}$ seconds). Under the same evaluation cost constraint, we compared (1) MLCF with Quasi-Monte Carlo points (QMC), (2) MLCF with Latin hypercube sampling (LHS), (3) MLCF with i.i.d points, (4) CF with i.i.d points, (5) MLMC with i.i.d points. The sample size is the optimal sample size for MLMC, which is listed in Table˜3. In this example, we used squared-exponential kernels. The hyper-parameters of the squared-exponential kernels are tuned independently at each level as illustrated in Appendix C.2. The kernel hyper-parameters at each level are tuned by maximizing the marginal likelihood associated with each $f_{l}-f_{l-1}$.
@@ D.4 Experimental Details for the Lotka-Volterra example
The Lotka-Volterra system [Lotka, 1925, 1927, Volterra, 1927] of ordinary differential equations is given by:
@@
for $t\in[0,s]$, with initial condition $u_{1}(0)=x_{5}$ and $u_{2}(0)=x_{6}$. To solve this system numerically, we use Stan [Carpenter et al., 2017]. Stan solves the ODE system for the times provided using the Dormand-Prince algorithm, a 4th/5th order Runge-Kutta method. For the multilevel construction, we consider a uniform time discretization with step size $h_{l}$ at level $l$. The grid points are $\{t_{i}=ih_{l}\}_{i=1}^{s/h_{l}}$. At higher levels $l$, smaller values of $h_{l}$ are chosen. $f_{l}$ is then defined as $f_{l}(\tilde{x})=(s)^{-1}h_{l}\sum^{s/h_{l}}_{i=1}u_{1}(t_{i})$, which corresponds to an approximation of the time average of the prey population $u_{1}(t)$ over the interval $[0,s]$.
@@ D.5 Experimental Details for Variational Inference of Bayesian Neural Networks
To demonstrate the effectiveness of the proposed method for variational inference, we utilize two-layer Bayesian neural networks, with the middle hidden layer having some hidden units, and use ReLU as the activation function. The prior of weights are set to be standard Gaussians $N(0,1)$. We use Adam as the stochastic optimizer to training the variational parameters $\lambda$ with initial learning rate $10^{-4}$ and a step-based decay function for the learning rate at iteration $l$: $\eta_{l}=\beta^{\text{floor}(l/r)}$ with drop parameter $r=250$ and $\beta=0.95$. For MC, MLMC and MLMCRG estimators, we use $5$ Monte Carlo samples to estimate the gradients during the optimization process while for MLCFRG estimators, we only use $1$ sample.
@@
We also study the effects of kernels used in variational inference for BNNs. In Figure˜9 and Figure˜9, we use square-exponential kernels for the cases when the number of hidden states is $15$ and $20$, respectively. The hyperparameters of kernels is chosen by median heuristics. For MC, MLMC and MLMCRG estimators, we use $5$ Monte Carlo samples to estimate the gradients during the optimization process while for MLCFRG estimators, we only use $1$ sample.
@@
We also study the effect of sample sizes: we test the effect of using the same strategy of sample sizes as in [Fujisawa and Sato, 2021]. We used preconditioned squared-exponential kernels in this example. The hyperparameters of kernels is chosen by median heuristics. For both Figure˜11 and Figure˜11, MLMCRG use the sample size strategy as in [Fujisawa and Sato, 2021] with starting sample size $n_{0}=5$. In Figure˜11, MLCFRG use the same sample size strategy as MLMCRG. In Figure˜11, sample size is fixed to be $1$ for MLCFRG for all iterations.

### v4

@@ 4 Experimental Results
To compare MCLF using quasi-Monte Carlo points (QMC), Latin hypercube sampling (LHS), and i.i.d points (IID) with MLMC, multilevel Bayesian quadrature (MLBQ) (Li et al., 2023) and CF using i.i.d points, we repeat the experiment 100 times. Multilevel methods in this example utilize the optimal sample size for MLMC. Details about the sample size, evaluation cost at each level, and other relevant information can be found in Appendix D.3. As shown in Figure˜4, under the same evaluation cost constraint, MLCF outperforms MLMC, MLBQ and CF. Figure˜4 also shows that experimental designs can improve the performance of MLCF.
@@
We now consider to perform Bayesian inference for the Lotka-Volterra system (Lotka, 1925; 1927; Volterra, 1927), which is also known as the predator-prey model. The model usually uses a system of differential equations:
@@
to describe the interaction between a predator and its prey in an ecosystem. $u_{1}(t)$ and $u_{2}(t)$ are the prey population and the predator population at time $t\in[0,s]$, for some $s\in\mathbb{R}_{+}$. The initial conditions of the system are $u_{1}(0)=x_{5}$ and $u_{2}(0)=x_{6}$. The observations $u_{1}(t_{i})$ and $u_{2}(t_{i})$ obtained exhibit log-normal noise with independent standard deviation $x_{7}$ and $x_{8}$ respectively, for all $i\in\{1,\ldots,m\}$. We can re-parameterize $x$ as in (Sun et al., 2023a; b) such that the re-parameterized model has parameters $\tilde{x}\in\mathbb{R}^{8}$. With the Gaussian distribution priors we assign on $\tilde{x}$ and the observations, we can construct the posterior distribution of $\tilde{x}$. The quantity of interest $\Pi[f]$ is the posterior expectation of the average prey population over the time period between 0 and $s$, i.e. $\Pi[f]=\int_{\mathbb{R}^{8}}f(\tilde{x})\pi(\tilde{x})d\tilde{x}$, where $\pi$ is the posterior probability distribution of $\tilde{x}$ and f(x~)f($\tilde${x}) is the average prey population between 0 and $s$ with the model parameter $\tilde{x}$. $f(\tilde{x})=s^{-1}\int_{0}^{s}u_{1}(t)dt$ is approximated with $(s)^{-1}h\sum^{s/h}_{i=1}u_{1}(t_{i})$, where $h$ is the step size and each $u_{1}(t_{i})$ is obtained by solving the differential equations numerically. The real-world dataset (Hewitt, 1921) consisting of the population of snowshoe hares (prey) and Canadian lynxes (predators) is used as observations for our study. With the real-world observations, we conduct Bayesian inference and use a MCMC sampler (no-U-turn sampler) in Stan (Carpenter et al., 2017) to obtain samples.
@@
We further extend the proposed MLCF estimators for variational inference, i.e. using MLCFRG estimators for the gradient of the ELBO. In particular, we applied a Bayesian neural network regression model to the UCI wine-quality-red dataset as (Fujisawa & Sato, 2021). The Bayesian neural network consists of a $15$-unit or a $20$-unit hidden layer with ReLU activations. For each weight $w_{i}$, we set its prior as Gaussian $w_{i}\sim N(\mu_{w},\sigma_{w})$, and the response variable $y\sim N(\phi(x,\{w_{i}\}_{i=1}^{d}),\sigma)$. The model then have a posterior of $d=392$ (when the number of hidden units is $15$) or $d=522$ (when the number of hidden units is $20$) and was applied to a dataset sub-sampled from the wine-quality-red dataset. The posterior is approximated by deploying a diagonal Gaussian distribution as the variational distribution. We compare the MLCFRG with MLMCRG, MLMC (with the number of levels being $2$) and MC estimators. For MLCFRG, we use preconditioned squared-exponential kernels (Oates et al., 2017); see details in Appendix C.1. The results, the average training ELBO/test log-likelihood (averaged over 10 runs) and associated standard deviations, are illustrated in Figure˜6 and Figure˜6. For each run, the networks are initialized under identical settings. See Appendix D.5 for extra details and experiments. Empirical results show that the MLCFRG estimators tend to outperform the other estimators as they have shown a faster convergence rate towards the optimal training ELBO and testing log-likelihood over number of iterations or equivalently the number of training data.
@@ 5 Conclusion
ACKNOWLEDGMENTS
Zhuo Sun is supported by Fundamental Research Funds for the Central Universities 2025110590 of Shanghai University of Finance and Economics.
Ethics Statement and Reproducibility Statement
This work raises no specific ethical concerns beyond standard practices in machine learning research. All methods, datasets, and hyperparameters are described in detail, and core code is released in supplementary materials.
@@ ACKNOWLEDGMENTS
The Use of Large Language Models (LLMs)
During the preparation of this manuscript, the authors used ChatGPT to polish the writing (e.g., improving grammar, readability, and clarity). The content, technical contributions, and conclusions of the paper were developed entirely by the authors, who take full responsibility for all ideas and results presented.
@@ A.2 Stein kernels
$k_{0}(x,x^{\prime})$ $=\nabla_{x}\cdot\nabla_{x^{\prime}}k(x,x^{\prime})+\nabla_{x}\log\pi(x)\cdot\nabla_{x^{\prime}}k(x,x^{\prime})+\nabla_{x^{\prime}}\log\pi(x^{\prime})\cdot\nabla_{x}k(x,x^{\prime})$
@@ Appendix B Proofs of Theoretical Results
due to the property of the Stein kernel $k^{l}_{0}$ that the Stein kernel $k^{l}_{0}$ satisfies $\int_{\mathcal{X}}k^{l}_{0}(x,x^{\prime})\pi(x)dx=0$ for all $x\in\mathcal{X}$ (Oates et al., 2017; 2019). Then, we have
@@
Following the proof of Theorem 1 of (Oates et al., 2019) or Theorem 11.13 of (Wendland, 2004), under assumptions A1-7, there exists $r^{*}_{l}>0$ and $h^{*}_{l}>0$, for $h_{l}<h^{*}_{l}$,
@@ C.1 Kernels and Their Derivatives
A preconditioned squared-exponential kernel (Oates et al., 2017) is:
@@ C.2 Hyper-parameter Selection
We propose to select kernel hyperparameters $\nu$ through a marginal likelihood objective by noticing the equivalence between the optimal Stein kernel-based control variates on the objectives in (Oates et al., 2017; Sun et al., 2023a) and the posterior mean of a zero-mean Gaussian process model with covariance matrix $k_{0}(x,y)$. See (Oates et al., 2017) for a discussion in the scalar-valued control variates case; see (Sun et al., 2023a) for a discussion in the vector-valued control variates case. Denotes all kernel-hyperparameters by $\nu:=\{\nu_{1},\ldots,\nu_{L}\}$, and we maximize the sum of the marginal likelihood,
@@ D.3 Experimental Details for the ODE example
We follow the settings described in (Giles, 2015; Li et al., 2023). For completeness, we provide details of the solver (finite difference approximation). The boundary-value ordinary differential equation (ODE) with random coefficient and random forcing is given by:
@@
For the ODE example, we have evaluation cost at each level $C=(C_{0},C_{1},C_{2})=(1.22,3.57,11.89)$ (all measured in $10^{-3}$ seconds). Under the same evaluation cost constraint, we compared (1) MLCF with Quasi-Monte Carlo points (QMC), (2) MLCF with Latin hypercube sampling (LHS), (3) MLCF with i.i.d points, (4) CF with i.i.d points, (5) MLMC with i.i.d points, (6) MLBQ with i.i.d points. The sample size is the optimal sample size for MLMC, which is listed in Table˜3. In this example, we used squared-exponential kernels. The hyper-parameters of the squared-exponential kernels are tuned independently at each level as illustrated in Appendix C.2. The kernel hyper-parameters at each level are tuned by maximizing the marginal likelihood associated with each $f_{l}-f_{l-1}$.
@@
Figure 8: Effect of Kernels and Kernel Hyper-parameters: Absolute integration error under a budget constraint (Y-axis log-scale). PSE stands for preconditioned square exponential kernels. SE stands for square exponential kernels. MH stands for median heuristic.
Extra Experiments: Effect of Kernels and Kernel Hyper-parameters
We also study the effect of kernels for the ODE example. We use i.i.d samples for all settings in Figure˜8. Under the same evaluation cost constraint, we compared (1) MLCF with the square exponential kernels (length-scale tuned by maximizing marginal likelihoods) (2) MLCF with the preconditioned square exponential kernels (length-scale tuned by maximizing marginal likelihoods) (3) MLCF with the square exponential kernels (length-scale selected by median heuristic) (4) MLMC (5) CF (length-scale tuned by maximizing marginal likelihoods). We still use the sample size listed in Table˜3. We find that setting the ‘length-scale’ of the square exponential kernels by median heuristic works well in practice.
@@ D.4 Experimental Details for the Lotka-Volterra example
The Lotka-Volterra system (Lotka, 1925; 1927; Volterra, 1927) of ordinary differential equations is given by:
@@
for $t\in[0,s]$, with initial condition $u_{1}(0)=x_{5}$ and $u_{2}(0)=x_{6}$. To solve this system numerically, we use Stan (Carpenter et al., 2017). Stan solves the ODE system for the times provided using the Dormand-Prince algorithm, a 4th/5th order Runge-Kutta method. For the multilevel construction, we consider a uniform time discretization with step size $h_{l}$ at level $l$. The grid points are $\{t_{i}=ih_{l}\}_{i=1}^{s/h_{l}}$. At higher levels $l$, smaller values of $h_{l}$ are chosen. $f_{l}$ is then defined as $f_{l}(\tilde{x})=(s)^{-1}h_{l}\sum^{s/h_{l}}_{i=1}u_{1}(t_{i})$, which corresponds to an approximation of the time average of the prey population $u_{1}(t)$ over the interval $[0,s]$.
@@ D.5 Experimental Details for Variational Inference of Bayesian Neural Networks
To demonstrate the effectiveness of the proposed method for variational inference, we utilize two-layer Bayesian neural networks, with the middle hidden layer having some hidden units, and use ReLU as the activation function. The prior of weights are set to be standard Gaussians $N(0,1)$. We use Adam as the stochastic optimizer to training the variational parameters $\lambda$ with initial learning rate $10^{-4}$ and a step-based decay function for the learning rate at iteration $l$: $\eta_{l}=\beta^{\text{floor}(l/r)}$ with drop parameter $r=250$ and $\beta=0.95$. For MC, MLMC and MLMCRG estimators, we use $5$ Monte Carlo samples to estimate the gradients during the optimization process while for MLCFRG estimators, we only use $1$ sample. The details of optimized training ELBO and test log-likelihood are presented in Table˜5 and Table˜6.
Table 5: Training ELBO and Test Log-likelihood (Hidden Dimension Size is $15$)
Method Training ELBO Test Log-likelihood
MC $6979.50~(\pm 1593.7)$ $1683.71~(\pm 542.1)$
MLMC $5214.09~(\pm 949.3)$ $1353.94~(\pm 395.8)$
MLMCRG $2187.77~(\pm 83.8)$ $388.42~(\pm 0.5)$
MLCFRG $2101.95~(\pm 36.3)$ $388.34~(\pm 0.9)$
Table 6: Training ELBO and Test Log-likelihood (Hidden Dimension Size is $20$)
Method Training ELBO Test Log-likelihood
MC $3581.99~(\pm 57.8)$ $771.04~(\pm 15.3)$
MLMC $3534.61~(\pm 63.9)$ $762.67~(\pm 27.5)$
MLMCRG $2010.09~(\pm 18.1)$ $386.27~(\pm 0.1)$
MLCFRG $2014.54~(\pm 20.2)$ $386.57~(\pm 0.3)$
@@
We also study the effects of kernels used in variational inference for BNNs. In Figure˜10 and Figure˜10, we use square-exponential kernels for the cases when the number of hidden states is $15$ and $20$, respectively. The hyperparameters of kernels is chosen by median heuristics. For MC, MLMC and MLMCRG estimators, we use $5$ Monte Carlo samples to estimate the gradients during the optimization process while for MLCFRG estimators, we only use $1$ sample.
@@
We also study the effect of sample sizes: we test the effect of using the same strategy of sample sizes as in (Fujisawa & Sato, 2021). We used preconditioned squared-exponential kernels in this example. The hyperparameters of kernels is chosen by median heuristics. For both Figure˜12 and Figure˜12, MLMCRG use the sample size strategy as in (Fujisawa & Sato, 2021) with starting sample size $n_{0}=5$. In Figure˜12, MLCFRG use the same sample size strategy as MLMCRG. In Figure˜12, sample size is fixed to be $1$ for MLCFRG for all iterations.